OpenStudy (anonymous):

Quadratic Equations: I just learned this yesterday and now understand it so I would like to help you guys. It took me like 3 - 4 days to learn.

OpenStudy (anonymous):

If you need help these are some vids I could tell you helps.

OpenStudy (anonymous):

https://www.youtube.com/watch?v=3ayhvAI3IeY

OpenStudy (danjs):

i like the enthusiasm. good work,

OpenStudy (anonymous):

https://www.youtube.com/watch?v=hn2zutY5UIU and https://www.youtube.com/watch?v=velArLW85MU these are both links for Quadratic Formula and the last tab I typed in is Factoring and Quadratic just to clarify

OpenStudy (anonymous):

that is all you need to look at if you need help

OpenStudy (anonymous):

the bottom two are filmed in a class and he makes jokes here and there if you get bored

OpenStudy (anonymous):

to remember the Quadratic Formula just remember this little skit. There was a negative boy who couldn't decide to go to a radical party but the boy was square and missed out on four awesome chicks. The party was over a 2am.

OpenStudy (danjs):

move from the general form for this a*x^2 + bx + c = 0 and solve for x, to derive the formula... it is sort of interesting

OpenStudy (anonymous):

that is the formula

OpenStudy (danjs):

yes show how you solve for x

OpenStudy (anonymous):

and if you need help with being able to fill it in you use this formula ax^2 + bx + c = 0

OpenStudy (anonymous):

lets just say you have x^2 + 4x - 5 = 0

OpenStudy (anonymous):

you would check to see if it is factorable

OpenStudy (danjs):

i mean, start with ax^2 + bx + c = 0 solve that for x,

OpenStudy (anonymous):

like what do you mean like explain it?

OpenStudy (danjs):

change the general form into the quadratic equation using algebra, start with the general form ax^2+bx+c=0 use algebra, solving for x will turn that into the familiar quadratic formula... it is pretty interesting the process to get there

OpenStudy (anonymous):

you mean solving the equation that have right there?

OpenStudy (anonymous):

or make an equation and then show how to solve it?

OpenStudy (mathmale):

Liam: Mind clarifying: Are you teaching this subject, asking questions about it, or what? It'd probably do you the most good to focus on one or the other, not both.

OpenStudy (anonymous):

does anyone know the codes of this site

OpenStudy (danjs):

i mean just solve ax^2 + bx + c = 0 , for x

OpenStudy (anonymous):

like the radicals how to type the equation

OpenStudy (danjs):

deriving the formula

OpenStudy (danjs):

i just use the equation button and the icons provided in that box

OpenStudy (anonymous):

lets just say you have x^2 + 5x - 11

OpenStudy (anonymous):

you would replace the -B with - 5x since it is asking for negative of b since the equation ax^2 + bx + c will tell you what a is b is and c is

OpenStudy (anonymous):

then in the radical you would replace b with 25 since it is b^2. Then you have -4ac which turns into 1 and -11 which you multiply and then you multiply -11 by -4 giving you 44. Now you have 25 + 44 there. Then now you have 2a. It becomes 2.

OpenStudy (anonymous):

my computer is bugging out so yea sorry for the delay

OpenStudy (anonymous):

b±√25+44 ____________ 2a

OpenStudy (anonymous):

I was looking up how to input those symbols

OpenStudy (anonymous):

the next step is adding the square root

OpenStudy (anonymous):

well the numbers in the radical

OpenStudy (anonymous):

which gives you √69 ______ 2

OpenStudy (anonymous):

what you do to the radical is find a way to divide it

OpenStudy (anonymous):

but since you can't divide this radical in to have an even number radical you can't solve it

OpenStudy (anonymous):

since there are always those equations that can't be solved

OpenStudy (anonymous):

with Quadratic Formula

OpenStudy (anonymous):

my laptop is bugging out

OpenStudy (anonymous):

my gosh I need to learn the code for this since I had to find the secondary laptop keyboard thing

OpenStudy (danjs):

sqrt ( ----) works for me

OpenStudy (mathmale):

Liam, my advice is that you concentrate on your own work for now. Once you can "ace" quadratic functions, quadratic equations and their graphs, THAT will be the time to offer to share your knowledge with others. Please, Liam, focus on specific math topics only if you choose to post questions here. We are not interested in the failure of your laptop. The Code of Conduct is here: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

sqrt(----)

OpenStudy (anonymous):

\[\sqrt(----)\]

OpenStudy (anonymous):

thx

OpenStudy (danjs):

you mean entering equations? hit the button under the input box for equation and then just use the buttons there for what you need

OpenStudy (danjs):

that latex language uses those squiggly brakets for a bunch of things, i just click the menu buttons mos the time

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I should try and explain another equation

OpenStudy (anonymous):

I just think I should use draw

OpenStudy (anonymous):

I didn't know it was closed lol

OpenStudy (anonymous):

I gtg

OpenStudy (mathmale):

Bye. Next time, be clear: Are you asking questions or are you explaining problems to others? Be certain you know which.